OpenStudy (anonymous):

the exact value of cos(- pi/4)

OpenStudy (anonymous):

1 over sqroot of 2

OpenStudy (anonymous):

 what would be the value of a,b,c

OpenStudy (callisto):

cos(- π/4) = cos (π /4) \[=1/\sqrt{2}\] Rationalise the answer \[=(1 \times\sqrt{2})/(\sqrt{2} \times \sqrt{2})\] Can you do it from here?

OpenStudy (anonymous):

not too sure

OpenStudy (anonymous):

.25

OpenStudy (anonymous):

how would you put it in the form i said above?

OpenStudy (anonymous):

actually it's 1! my B! it's actully 0.99990604980155050801971795294044 to be exact! concure?

OpenStudy (anonymous):

what the value for a,b,c

OpenStudy (anonymous):

is it 1 for a, 2 for b, and 2 for c

OpenStudy (callisto):

Continue with my workings: \[= \sqrt{2} /2\] Can you get the answer now?

OpenStudy (anonymous):

listen to callisto!

OpenStudy (anonymous):

i know the answer but i want to know that the values are correct. cause the question only asks for the excat value of the 3 letters in the drawing above!

OpenStudy (anonymous):

I got the answer. Thanks guys!

OpenStudy (callisto):

Sorry, i thought you can work it out on your own..